# Boosting Transferability in Vision-Language Attacks via Diversification along the Intersection Region of Adversarial Trajectory

### v1

@@ Abstract
Vision-language pre-training (VLP) models exhibit remarkable capabilities in comprehending both images and text, yet they remain susceptible to multimodal adversarial examples (AEs). Strengthening adversarial attacks and uncovering vulnerabilities, especially common issues in VLP models (e.g., high transferable AEs), can stimulate further research on constructing reliable and practical VLP models. A recent work (i.e., Set-level guidance attack) indicates that augmenting image-text pairs to increase AE diversity along the optimization path enhances the transferability of adversarial examples significantly. However, this approach predominantly emphasizes diversity around the online adversarial examples (i.e., AEs in the optimization period), leading to the risk of overfitting the victim model and affecting the transferability. In this study, we posit that the diversity of adversarial examples towards the clean input and online AEs are both pivotal for enhancing transferability across VLP models. Consequently, we propose using diversification along the intersection region of adversarial trajectory to expand the diversity of AEs. To fully leverage the interaction between modalities, we introduce text-guided adversarial example selection during optimization. Furthermore, to further mitigate the potential overfitting, we direct the adversarial text deviating from the last intersection region along the optimization path, rather than adversarial images as in existing methods. Extensive experiments affirm the effectiveness of our method in improving transferability across various VLP models and downstream vision-and-language tasks (e.g., Image-Text Retrieval(ITR), Visual Grounding(VG), Image Captioning(IC)).
@@
Vision-language pre-training (VLP) models utilize multimodal learning, leveraging large-scale image-text pairs to bridge the gap between visual and language understanding. These models showcase revolutionary performance across various downstream Vision-and-Language tasks, including image-text retrieval, image captioning, visual grounding, and visual entailment, as demonstrated in [18, 14, 19, 34]. Despite their success, recent research highlights the high vulnerability of VLP models, particularly when confronted with multimodal adversarial examples [47, 27]. Uncovering the vulnerabilities, especially common issues, can drive further research on building reliable and practical VLP models.
@@ 2.1 Vision-Language Pre-training Models
There are two popular approaches for VLP models in learning vision-language representations: the fused architecture and the aligned architecture. Fused VLP models (e.g., ALBEF [21], TCL [43]), initially employ two separate unimodal encoders to learn features for text and images. Subsequently, a multimodal encoder is utilized to fuse the embeddings of text and images. In contrast, aligned VLP models, exemplified by CLIP [32], focus on aligning the feature spaces of distinct unimodal encoders. This paper concentrates on evaluating our proposed method using multiple popular fused and aligned VLP models.
@@ 2.3 Transferability of Adversarial Examples
Adversarial attacks [16, 12, 11] are typically categorized as white-box and black-box attacks. In a white-box setting [17, 15], the attacker has full access to the model, whereas black-box attacks [3, 30], more realistic in practical applications, occur when information about the model is limited. In the realm of image attacks, prevalent methods for crafting transferable adversarial examples often leverage data augmentation techniques (e.g., DIM [42], TIM [8], SIM[25], ADMIX [40], PAM [48]). Zhang et al. [47] introduced a white-box attack targeting popular VLP models for downstream tasks in the multimodal domain. Building upon this work, Lu et al. [27] proposed SGA, considering the diversity of adversarial examples by expanding single image-text pairs to sets of images and texts to conduct black-box attacks on VLP models.
@@ 3.1 Background and Motivation
| ALBEF | TCL | CLIP | CLIP | ||||||
@@
To further examine the impact of image augmentation along the optimization path in the SGA method, we utilize ALBEF as a surrogate model to generate multimodal adversarial examples. These examples are then employed to target VLP models such as TCL and CLIP, assessing the transferability of the attacks. Detailed results are presented in Table 1. Our observations reveal that SGA enhances the transferability of adversarial attacks, showing an increase ranging from 6.14% to 17.81%. However, it is noteworthy that the success rate of attacks on the target models remains notably lower than that on the source model. This discrepancy is primarily attributed to the fact that SGA predominantly emphasizes diversity around adversarial examples during the optimization period, without adequately considering the diversity of adversarial examples toward the clean image, leading to the risk of overfitting the victim model and affecting the transferability.
@@ 3.3 Text-guided Augmentation Selection
3.4 Adversarial Texts deviating from the Intersection Region
@@ 4.1 Setups
VLP Models. In our transferability evaluation experiments across various VLP models, we explore two typical architectures: fused and aligned VLP models. We select CLIP [32] for the aligned VLP model. CLIP offers a choice between two distinct image encoders, namely ViT-B/16 [9] and ResNet-101 [13], denoted as CLIP and CLIP, respectively. In the case of fused VLP models, we opt for ALBEF [21] and TCL [43]. ALBEF’s image encoder is a 12-layer ViT-B/16 visual transformer, complemented by two 6-layer transformers for the text encoder and multimodal encoder. TCL shares the same model architecture as ALBEF but differs in its pre-training objectives, maintaining a common fundamental model structure.
@@
Adversarial Attack Settings. In our study, we adopt adversarial attack settings of SGA [27] to ensure a fair comparison. Specially, we leverage BERT-Attack [22] to craft adversarial texts. The perturbation bound is set as 1 (See Figure 2) and length of word list . PGD [28] is employed to get adversarial images and the perturbation bound, denoted as , is set as 8/255. Additionally, iteration steps is set as 10 and each step size . Furthermore, when randomly sampling from the intersection region of adversarial trajectory, we set the number of samples to 5. When generating adversarial texts, we set the three parameters in Equation 6 to 0.6, 0.2, and 0.2 respectively. The values chosen for these adjustable parameters can be found in Section 4.2.
Evaluation Metrics. The primary evaluation metric for assessing adversarial transferability is the Attack Success Rate (ASR). ASR represents the percentage of successful attacks among the total number of adversarial examples generated. A higher ASR indicates increased effectiveness and transferability of the adversarial attack.
@@ 4.2 Optimal Parameters
| ALBEF | TCL | CLIP | CLIP | ||||||
@@ 4.3 Cross-Model Transferability
| ALBEF | TCL | CLIP | CLIP | ||||||
@@
As outlined in Section 4.1, our experimental design focuses on assessing the transferability of adversarial examples across two widely adopted VLP model architectures: fused and aligned. There are four VLP models selected, namely: ALBEF, TCL, CLIP, CLIP. The selected downstream V+L task for evaluation is image-text retrieval. Our approach involves employing one of four models to generate multimodal adversarial examples within the specified parameters of our proposed method. Subsequently, we validate the effectiveness of the generated adversarial examples through a comprehensive set of experiments, encompassing both self-attacks and attacks on the other three models. This evaluation encompasses one white-box attack and three distinct black-box attacks.
@@
Cross-Model Transferability in Same Architecture. ALBEF and TCL both utilize a similar model architecture but differ in their pre-training objectives while maintaining a common fundamental model structure. Therefore, when ALBEF and TCL are employed as target models for each other, the success rate of attacks using multimodal adversarial examples is remarkably high. Notably, the black-box attack success rate of IR reaches 95.58% when TCL is the target model in our proposed method. Additionally, the transferability of SGA can reach approximately 90%, but the room for improvement is very limited, our method has improved compared to SGA, ranging from 2.71% to 3.69%. In contrast, when considering CLIP and CLIP—both being aligned VLP models—their image encoders vary significantly, with one utilizing the Vision Transformer and the other employing ResNet-101. Given the substantial structural differences between traditional CNN networks and Transformers, the existing methods have relatively low cross-model transferability between them, indicating significant room for improvement. Our method outperforms SGA by approximately 7.67% to 11.62%.
@@ 4.4 Cross-Task Transferability
We not only assess the transferability of multimodal adversarial examples generated by our proposed method across different VLP models but also conduct experiments to evaluate its effectiveness in transferring across diverse V+L tasks. Specifically, we craft adversarial examples for the Image-Text Retrieval (ITR) task and evaluate them on Visual Grounding (VG) and Image Captioning (IC) tasks. As evident from Table 4 and visual results in Figure 3 and 4, the adversarial examples generated for ITR demonstrate transferability, successfully impacting both VG and IC tasks. This highlights the efficacy of cross-task transferability in our proposed method. Furthermore, our transferability consistently outperforms that of SGA.
@@ 5 Conclusion
Appendix 0.A Intersection Regional Optimal Sample
@@
To achieve the aforementioned objectives, we employ ALBEF [21], TCL [43], and CLIP [32] as surrogate models on the Flickr30K [31] dataset to generate multimodal adversarial examples, respectively. The average value is computed across all examples. Given that we set the PGD [28] step size to 10 and utilize random noise initialization, there will be 9 iterations of sampling and selecting the optimal sample. The detailed results are illustrated in Figure 6.
First, let’s analyze the numerical distribution of the three ratios: , , and . As depicted in Figure 6, it is evident that, for any VLP model and at each optimization step, the proportion of adversarial images consistently remains the highest, typically surpassing 60%. Furthermore, in the optimal samples, the proportion of adversarial images does not exceed 85%. In the case of SGA, it can be viewed as a special scenario within our method, wherein the proportion of adversarial images in the selected samples consistently reaches 100% in each iteration.
@@
Next, we delve into an examination of the dynamic shifts in the optimal sample proportions occupied by adversarial images, historical adversarial images, and clean images throughout the iteration process. The specific outcomes are depicted in Figure 6. Across the four VLP models, a discernible diminishing trend is observed in the proportion occupied by adversarial images. In the PGD multi-step iterative attack method, a single-step attack undergoes refinement through the incorporation of multiple successive perturbations. In the initial iterations, owing to the relatively modest accumulation of perturbations in the adversarial images, these perturbations fall below the constraint limits. Consequently, during these early steps, the optimal samples tend to align closely with the adversarial images. As the perturbations in the adversarial images continue to accrue, they eventually surpass the imposed constraints. The portions exceeding these constraints are subsequently adjusted to the constraint value, e.g., 8/255. At this juncture, an excessive reliance on adversarial images inevitably leads to a decline in the effectiveness of the attack. However, judicious consideration of the adversarial perturbations around clean images and the most recent adversarial image helps mitigate the risk of overfitting. Furthermore, our analysis extends to the changing trend in the proportion of clean images within the optimal samples. Notably, for the CLIP model, there is a conspicuous upward pattern, while the other three VLP models also exhibit an upward trend, albeit at a very slow pace. Across all VLP models, the proportion of the last adversarial image shows a slow increase over the first 7 or 8 steps, followed by a decrease in the final two or three steps.
Appendix 0.B Adversarial Transferability on Large Language Models
@@
Recently, Large Language Models (LLMs) have garnered widespread applications. Their utilization spans various domains, showcasing their versatility and efficacy in tackling complex tasks. To further explore the adversarial robustness of LLMs, we conduct specialized extension experiments. Specifically, we utilize ALBEF as a surrogate model, generating adversarial examples under settings with perturbation constraints of 16/255 and single-step perturbation of 0.5/255, with a step size of 500. Furthermore, we evaluate them on advanced LLMs (e.g., GPT-4 [1], Claude-3) by prompting these systems with the query "Describe this image". The results, depicted in Figure 7 and 8, demonstrate that adversarial examples generated by our method can mislead today’s state-of-the-art LLMs into producing incorrect answers.
Appendix 0.C Detailed Results about Cross Model Transferability

### v3

@@ Abstract
Vision-language pre-training (VLP) models exhibit remarkable capabilities in comprehending both images and text, yet they remain susceptible to multimodal adversarial examples (AEs). Strengthening attacks and uncovering vulnerabilities, especially common issues in VLP models (e.g., high transferable AEs), can advance reliable and practical VLP models. A recent work (i.e., Set-level guidance attack) indicates that augmenting image-text pairs to increase AE diversity along the optimization path enhances the transferability of adversarial examples significantly. However, this approach predominantly emphasizes diversity around the online adversarial examples (i.e., AEs in the optimization period), leading to the risk of overfitting the victim model and affecting the transferability. In this study, we posit that the diversity of adversarial examples towards the clean input and online AEs are both pivotal for enhancing transferability across VLP models. Consequently, we propose using diversification along the intersection region of adversarial trajectory to expand the diversity of AEs. To fully leverage the interaction between modalities, we introduce text-guided adversarial example selection during optimization. Furthermore, to further mitigate the potential overfitting, we direct the adversarial text deviating from the last intersection region along the optimization path, rather than adversarial images as in existing methods. Extensive experiments affirm the effectiveness of our method in improving transferability across various VLP models and downstream vision-and-language tasks. Code is available at https://github.com/SensenGao/VLPTransferAttack.
@@
Vision-language pre-training (VLP) models utilize multimodal learning, leveraging large-scale image-text pairs to bridge the gap between visual and language understanding. These models showcase revolutionary performance across various downstream Vision-and-Language tasks, including image-text retrieval, image captioning, visual grounding, and visual entailment, as demonstrated in [28, 21, 29, 45]. Despite their success, recent research underscores the significant vulnerability of VLP models, particularly when confronted with multimodal adversarial examples [59, 37, 17, 18, 8, 38, 13]. Uncovering the vulnerabilities, especially common issues, can drive further research aimed at building more reliable and practical VLP models.
@@ 2.1 Vision-Language Pre-training Models
There are two popular approaches for VLP models in learning vision-language representations: the fused architecture and the aligned architecture. Fused VLP models (e.g., ALBEF [31], TCL [55]), initially employ two separate unimodal encoders to learn features for text and images. Subsequently, a multimodal encoder is utilized to fuse the embeddings of text and images. In contrast, aligned VLP models, exemplified by CLIP [43], focus on aligning the feature spaces of distinct unimodal encoders and benefit downstream tasks significantly [1]. This paper concentrates on evaluating our proposed method using multiple popular fused and aligned VLP models.
@@ 2.3 Transferability of Adversarial Examples
Adversarial attacks [26, 16, 15, 19, 14] are typically categorized as white-box and black-box attacks. In a white-box setting [27, 24], the attacker has full access to the model, whereas black-box attacks [4, 41], more realistic in practical applications, occur when information about the model is limited. In the realm of image attacks [23, 25, 22], prevalent methods for crafting transferable adversarial examples often leverage data augmentation techniques (e.g., DIM [54], TIM [10], SIM[35], ADMIX [51], PAM [60]). Zhang et al. [59] introduced a white-box attack targeting popular VLP models for downstream tasks in the multimodal domain. Building upon this work, Lu et al. [37] proposed SGA, considering the diversity of adversarial examples by expanding single image-text pairs to sets of images and texts to conduct black-box attacks on VLP models.
@@ 3.1 Background and Motivation
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||
@@
To further examine the impact of image augmentation along the optimization path in the SGA method, we utilize ALBEF as a surrogate model to generate multimodal adversarial examples. These examples are then employed to target VLP models such as TCL and CLIP, assessing the transferability of the attacks. Detailed results are presented in Table 1. Our observations reveal that SGA enhances the transferability of adversarial attacks, showing an increase ranging from 6.14% to 17.81%. However, it is noteworthy that the success rate of attacks on the target models remains notably lower than that on the source model. This discrepancy is primarily attributed to the fact that SGA predominantly emphasizes diversity around AE during the optimization period, without adequately considering the diversity of adversarial examples toward the clean image, leading to the risk of overfitting the victim model and affecting the transferability.
@@ 3.3 Text-guided Augmentation Selection
3.4 Adversarial Text deviating from the Intersection Region
@@ 4.1 Setups
VLP Models. In our transferability evaluation experiments across various VLP models, we explore two typical architectures: fused and aligned VLP models. We select CLIP [43] for the aligned VLP model. CLIP offers a choice between two distinct image encoders, namely ViT-B/16 [11] and ResNet-101 [20], denoted as CLIPViT and CLIPCNN, respectively. In the case of fused VLP models, we opt for ALBEF [31] and TCL [55]. ALBEF uses a 12-layer ViT-B/16 image encoder and two 6-layer transformers for text and multimodal encoding. TCL shares this architecture but has different pre-training objectives.
@@
Adversarial Attack Settings. In our study, we adopt adversarial attack settings of SGA [37] to ensure a fair comparison. Specifically, we leverage BERT-Attack [32] to craft adversarial texts. The perturbation bound is set as 1 and length of word list . PGD [39] is employed to get adversarial images and the perturbation bound, denoted as , is set as 8/255. Additionally, iteration steps is set as 10 and each step size . Furthermore, when randomly sampling from the intersection region of adversarial trajectory, we set the number of samples to 5. When generating adversarial texts, we set the three parameters in Equation 6 to 0.6, 0.2, and 0.2 respectively. The values chosen for these adjustable parameters can be found in Section 4.2.
Evaluation Metrics. The key metric for adversarial transferability is the Attack Success Rate (ASR), which measures the percentage of successful attacks among all generated adversarial examples. A higher ASR indicates more effective and transferable attacks.
@@ 4.2 Optimal Parameters
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||
@@ 4.3 Cross-Model Transferability
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||
@@
As outlined in Section 4.1, our experimental design focuses on assessing the transferability of adversarial examples across two widely adopted VLP model architectures: fused and aligned. There are four VLP models selected, namely: ALBEF, TCL, CLIPViT, CLIPCNN. The selected downstream V+L task for evaluation is image-text retrieval. Our approach involves employing one of four models to generate multimodal adversarial examples within the specified parameters of our proposed method. Subsequently, we validate the effectiveness of the generated adversarial examples through a comprehensive set of experiments, encompassing both self-attacks and attacks on the other three models. This evaluation encompasses one white-box attack and three distinct black-box attacks.
@@
Cross-Model Transferability in Same Architecture. ALBEF and TCL both utilize a similar model architecture but differ in their pre-training objectives while maintaining a common fundamental model structure. Therefore, when ALBEF and TCL are employed as target models for each other, the success rate of attacks using multimodal adversarial examples is remarkably high. Notably, the black-box attack success rate of IR reaches 95.58% when TCL is the target model in our proposed method. Additionally, the transferability of SGA can reach approximately 90%, but the room for improvement is very limited, our method has improved compared to SGA, ranging from 2.71% to 3.69%. In contrast, when considering CLIPViT and CLIPCNN—both being aligned VLP models—their image encoders vary significantly, with one utilizing the Vision Transformer and the other employing ResNet-101. Given the significant structural differences between traditional CNNs and Transformers, existing methods show low cross-model transferability, highlighting room for improvement. Our method outperforms SGA by about 7.67% to 11.62%.
@@ 4.4 Cross-Task Transferability
We not only assess the transferability of multimodal adversarial examples generated by our proposed method across different VLP models but also conduct experiments to evaluate its effectiveness in transferring across diverse V+L tasks. Specifically, we craft adversarial examples for the Image-Text Retrieval (ITR) task and evaluate them on Visual Grounding (VG) and Image Captioning (IC) tasks. As evident from Table 4 and visual results in Figure 2 and 3, the adversarial examples generated for ITR demonstrate transferability, successfully impacting both VG and IC tasks. This highlights the efficacy of cross-task transferability in our proposed method. Furthermore, our transferability consistently outperforms that of SGA.
@@ 5 Conclusion
Acknowledgments
This research is supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), Career Development Fund (CDF) of Agency for Science, Technology and Research (A*STAR) (No.: C233312028), and National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative (No. DTC-RGC-04).
Appendix 0.A Visualization on Multimodal Dataset
In Figure A1, we present the adversarial perturbations of our method for both image and text modalities. As mentioned in our experimental setup, we perturb only one token for text and use an 8/255 perturbation for images. The first two lines represent the clean image and the adversarial image, respectively. The third line indicates the added adversarial perturbation. The last two lines represent the clean text and the adversarial text, where green highlights the original words, while red highlights the words modified in the adversarial text.
Appendix 0.B Attack Formulation
In Algorithm 1, we provide a detailed demonstration of the process by which our method generates adversarial images and adversarial texts.
Appendix 0.C Intersection Regional Optimal Sample
@@
To achieve the aforementioned objectives, we employ ALBEF [31], TCL [55], and CLIP [43] as surrogate models on the Flickr30K [42] dataset to generate multimodal adversarial examples, respectively. The average value is computed across all examples. Given that we set the PGD [39] step size to 10 and utilize random noise initialization, there will be 9 iterations of sampling and selecting the optimal sample. The detailed results are illustrated in Figure A2.
First, let’s analyze the numerical distribution of the three ratios: , , and . As depicted in Figure A2, it is evident that, for any VLP model and at each step, the proportion of adversarial images consistently remains the highest, typically surpassing 60%. Furthermore, in the optimal samples, the proportion of adversarial images does not exceed 85%. In the case of SGA, it can be viewed as a special scenario within our method, wherein the proportion of adversarial images in the selected samples consistently reaches 100% in each iteration.
@@
Next, we delve into an examination of the dynamic shifts in the optimal sample proportions occupied by adversarial images, historical adversarial images, and clean images throughout the iteration process. The specific outcomes are depicted in Figure A2. Across the four VLP models, a discernible diminishing trend is observed in the proportion occupied by adversarial images. In the PGD multi-step iterative attack method, a single-step attack undergoes refinement through the incorporation of multiple successive perturbations. In the initial iterations, owing to the relatively modest accumulation of perturbations in the adversarial images, these perturbations fall below the constraint limits. Consequently, during these early steps, the optimal samples tend to align closely with the adversarial images. As the perturbations in the adversarial images continue to accrue, they eventually surpass the imposed constraints. The portions exceeding these constraints are subsequently adjusted to the constraint value, e.g., 8/255. At this juncture, an excessive reliance on adversarial images inevitably leads to a decline in the effectiveness of the attack. However, judicious consideration of the adversarial perturbations around clean images and the most recent adversarial image helps mitigate the risk of overfitting. Furthermore, our analysis extends to the changing trend in the proportion of clean images within the optimal samples. Notably, for the CLIPCNN model, there is a conspicuous upward pattern, while the other three VLP models also exhibit an upward trend, albeit at a very slow pace. Across all VLP models, the proportion of the last adversarial image shows a slow increase over the first 7 or 8 steps, followed by a decrease in the final two or three steps.
Appendix 0.D Rationale behind transferability improvement
The main reason for the low transferability of an adversarial example(AE) is that it heavily relies on the surrogate vision-language model, also known as the overfitting issue.
Why is SGA effective? SGA conducts augmentations around the online AEs (See Figure 1 (a)) to improve transferability “instead of increasing the overfitting risk". However, the augmentations still highly rely on the examples along the adversarial trajectory, leading to overfitting risks.
Why is our method effective? Unlike SGA, our method performs augmentations around online AEs considering the raw clean example and the AE at the previous step (See Figure 1 (b)). We utilize the clean example since it is entirely independent of the surrogate model and offers the potential to mitigate the risk of overfitting.
Overfitting analysis of SGA & our method via IIWs. To quantify overfitting risk, we use the PAC-Bayes theorem to measure the amount of information stored in the network’s weights (IIW)[52], a promising indicator of generalization ability. Smaller IIW means lower overfitting risks. Given an AE from SGA or our method during optimization iteration, we calculate its IIWs by feeding it to four VLMs, respectively. We count IIWs of 1000 AEs during optimization and show the averaging results in Figure A5. During optimization, the IIW of AEs from the SOTA baseline (i.e., SGA) initially decreases before sharply increasing. In contrast, our method leads to a continuous low IIW value in generated AEs, demonstrating that our method can overcome the overfitting risks. As a result, the attack transferability can be enhanced.
Appendix 0.E Adversarial Transferability on Large Language Models
@@
Recently, Large Language Models (LLMs) have garnered widespread applications. Their utilization spans various domains, showcasing their versatility and efficacy in tackling complex tasks. To further explore the adversarial robustness of LLMs, we conduct specialized extension experiments. Specifically, we utilize ALBEF as a surrogate model, generating adversarial examples under settings with perturbation constraints of 16/255 and single-step perturbation of 0.5/255, with a step size of 500. Furthermore, we evaluate them on advanced LLMs (e.g., GPT-4 [2], Claude-3) by prompting these systems with the query "Describe this image". The results, depicted in Figure A3 and A4, demonstrate that adversarial examples generated by our method can mislead today’s state-of-the-art LLMs into producing incorrect answers.
We have shown two cases on open VLP large models above. Here, we further extend experiments by randomly selecting 100 images from the Flick30K and generating adversarial images using SGA and our method with ALBEF as the surrogate, under the same perturbation setting (16/255). Then, we feed adversarial images into LLMs and query "Describe this image", and rank descriptions with 100 irrelevant texts based on CLIP similarity. If the description doesn’t rank top-N, consider a successful R@N attack. Table A1 illustrates ASR(%), showing our superior transferability to LLMs compared to SGA.
| GPT-4 | Claude-3 | Qwen-VL | |||||||
| Attack | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 |
| No Attack | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| SGA | 6.0 | 0.0 | 0.0 | 18.0 | 4.0 | 3.0 | 6.0 | 2.0 | 0.0 |
| Ours | 16.0 | 6.0 | 4.0 | 22.0 | 12.0 | 9.0 | 15.0 | 6.0 | 1.0 |
Appendix 0.F Time cost comparison.
Table A2 compares the time costs per AE between SGA and our method. and respectively represent the time for generating adversarial images and text. Our method takes 1.58-2.28 times longer overall compared to SGA, with generating adversarial images taking 2.33 to 2.59 times as long. This additional time primarily arises from the process of selecting optimal samples in the intersection region, involving a sample number of additional forward/backward passes compared to SGA. In the future, we will explore methods to select optimal samples to reduce computational costs.
| ALBEF | TCL | CLIPViT | CLIPCNN | |||||
| Attack | ||||||||
| SGA | 2.02 | 1.66 | 1.99 | 1.64 | 1.02 | 0.56 | 0.78 | 0.36 |
| Ours | 4.60 | 4.27 | 4.53 | 4.21 | 1.85 | 1.45 | 1.23 | 0.84 |
Appendix 0.G Detailed Results about Cross Model Transferability